Yield to maturity 6 years left to maturity. Interest is paid annually, and the bonds have a $1,000 par value and a coupon rate of 10%. a. What is the yield to maturity at a current market price of 1. $818? Round your answer to two decimal places. 2. $1,205? Round your answer to two decimal places.
Answer #1

Answer 1

Yield To Maturity(YTM) = (interest per period+ ((Redemption price - Current market price) / life remaining to maturity)) / ((.4*Redemption price)+ (.6*Current market price))

= ((1000*10%)+((1000-818)/6))/(.4*1000+.6*818)

= (100+30.3333333333)/890.8

= 130.333333333/890.8

= 0.14631043257

= 14.63%

Answer 2

Yield To Maturity(YTM) = (interest per period+ ((Redemption price - Current market price) / life remaining to maturity)) / ((.4*Redemption price)+ (.6*Current market price))

= ((1000*10%)+((1000-1205)/6))/(.4*1000+.6*1205)

= (100-34.1666666667)/1123

= 65.8333333333/1123

= 0.05862273671

= 5.86%